OpenStudy (anonymous):

The length of a rectangle is 5 ft more than twice its width, and the area of the rectangle is 63ft^2 . Find the dimensions of the rectangle. Width= Length=

OpenStudy (saifoo.khan):

Let the width = X. The length is twice the width plus 5. So the length must be 2X + 5. Area = 63 ft squared X multiplied by 2X + 5 2X squared + 25. 2X squared = 63 - 25 38 2X square root of 38. Therefore the width is the square root of 38 divided by 2 and the length is the square root of 38 plus 5.

OpenStudy (anonymous):

Ahh I was JUST about to say nevermind, I figured it out :D Thank you though man

OpenStudy (saifoo.khan):

Yuw. :)